A rod of length L has its ends A and B maintained at 20 °C and 40 °C respectively
until steady state condition prevail. The temperature at A is suddenly raised to
50 °C while that at B is lowered to 10 °C and maintained thereafter. Find the
subsequent temperature distribution of the rod.
1
Expert's answer
2020-12-23T17:07:40-0500

The heat equation:

"\frac{\partial u}{\partial t}=\alpha^2\frac{\partial^2u}{\partial x^2}"

In steady state:

"\frac{\partial^2u}{\partial x^2}=0"

Solving, we get

"u=ax+b"

The initial conditions, in steady–state, are 

"u=20, x=0"

"u=40, x=L"

Then:

"b=20, a=20\/L"

The temperature function in steady–state is

"u(x)=20x\/L+20"


The boundary conditions in the transient state are 

"u(0,t)=50, u(L,t)=10, u(x,0)=20x\/L+20"


We break up the required funciton u (x,t) into two parts:

"u(x,t)=u_s(x)+u_t(x,t)"

"u_s(x)" is a steady state solution, "u_t(x,t)" is a transient solution


"u_s(x)=ax+b"

"b=50, a=-40\/L"

"u_s(x)=-40x\/L+50"


"u_t(x,t)=u(x,t)-u_s(x)"

Putting boundary conditions, we have the boundary conditions relative to the transient solution :

"u_t(0,t)=u(0,t)-u_s(0)=50-50=0"

"u_t(L,t)=u(L,t)-u_s(L)=10-10=0"

"u_t(x,0)=u(x,0)-u_s(x)=20x\/L+20+40x\/L-50=60x\/L-30"


We have:

"u_t(x,t)=(Acos\lambda x+Bsin\lambda x)e^{-\alpha^2\lambda^2t}"

Using boundary conditions, we get:

"A=0, \lambda=\pi n/L"

Then:

"u_t(x,t)=Bsin\frac{\pi nx}{L}e^{-\frac{\alpha^2\pi^2 n^2}{L^2}t}"

The most general solution:

"u_t(x,t)=\displaystyle\sum_{n=1}^{\infin}B_nsin\frac{\pi nx}{L}e^{-\frac{\alpha^2\pi^2 n^2}{L^2}t}"

Using boundary conditions:

"u_t(x,0)=\displaystyle\sum_{n=1}^{\infin}B_nsin\frac{\pi nx}{L}=60x/L-30, 0<x<L"

"B_n=\frac{2}{L}\displaystyle\intop_0^L(60x/L-30)sin\frac{\pi nx}{L}dx="

"=\frac{2}{L}((60x/L-30)\frac{-cos(\pi nx/L)}{\pi n/L}-\frac{60}{L}\cdot\frac{-sin(\pi nx/L)}{\pi^2n^2/L^2})|^L_0="

"=\frac{2}{L}(-\frac{30Lcos(\pi n)}{\pi n}-\frac{30L}{\pi n})=-\frac{60}{\pi n}(cos(\pi n)+1)"


"B_n=-\frac{60(1+(-1)^n)}{\pi n}"


"B_n=0" when "n" is odd


"B_n=-\frac{120}{\pi n}" when "n" is even


"u_t(x,t)=\displaystyle\sum_{n=2,4,6,..}^{\infin}(-\frac{120}{\pi n}sin\frac{\pi nx}{L})e^{-\frac{\alpha^2\pi^2n^2}{L^2}t}"


"u(x,t)=-40x/L+50+\displaystyle\sum_{n=2,4,6,..}^{\infin}(-\frac{120}{\pi n}sin\frac{\pi nx}{L})e^{-\frac{\alpha^2\pi^2n^2}{L^2}t}"